# Multi-Agent Code Verification with Compound Vulnerability Detection

### v1

@@ 7.7 Impact
Risks: Over-reliance could reduce human review, missing novel bugs. The 50% FPR causes alert fatigue without good UX. Orgs might think 76% TPR means “catches all bugs”—24% false negative rate means human oversight essential. Compound detection relies on predefined patterns, missing emerging exploits.
@@ 8 Conclusion
LLMs generate buggy code: 29.6% of SWE-bench patches fail, 62% of BaxBench solutions have vulnerabilities. We built CodeX-Verify, a multi-agent system with mathematical foundations and compound vulnerability detection, addressing the 40–60% bug rate in LLM code.
@@
Testing on 300 Claude Sonnet 4.5 patches runs at ¡200ms per sample, flagging 72% for correction with 100% compound vulnerability detection. Our 50% FPR exceeds test-based methods (8.6%) because we flag security and quality issues that tests miss, a tradeoff for enterprise security.
[…]
Three directions: (1) Hybrid static-dynamic verification combining our framework with test execution for comprehensive coverage and low false positives. (2) Expanding from 4 to 100+ attack chains using security databases. (3) Multi-language support (C/C++, Java, TypeScript) via tree-sitter.

### v3

@@ 7.6 Impact
Risks: Over-reliance could reduce human review, missing novel bugs. The 50% FPR causes alert fatigue without good UX. Orgs might think 76% TPR means “catches all bugs”—24% false negative rate means human oversight essential.
@@ 8 Conclusion
LLMs generate buggy code: 29.6% of SWE-bench patches fail, 62% of BaxBench solutions have vulnerabilities. We built CodeX-Verify, a multi-agent verification system with rigorous information-theoretic foundations, addressing the 40–60% bug rate in LLM code.
We proved via submodularity of mutual information that combining agents with conditionally independent detection patterns finds more bugs than any single agent: $I(A_{1},\ldots,A_{n};B)>\max_{i}I(A_{i};B)$ when multiple agents are informative. Measured agent correlation $\rho=0.05$$0.25$ shows low redundancy across agents. We also proved diminishing returns (marginal gains decrease monotonically) via submodularity, confirmed by measured gains of +14.9pp, +13.5pp, +11.2pp.
Testing on 300 Claude Sonnet 4.5 patches runs at ¡200ms per sample, flagging 72% for correction. Our 50% FPR exceeds test-based methods (8.6%) because we flag security and quality issues that tests miss, a tradeoff for enterprise security.
@@
Three directions: (1) Hybrid static-dynamic verification combining our framework with test execution for comprehensive coverage and low false positives. (2) Multi-language support (C/C++, Java, TypeScript) via tree-sitter’s unified AST interface. (3) Learned threshold optimization on larger datasets to reduce FPR while maintaining TPR.